OpenStudy (anonymous):

what are both solutions for the quadratic equation below: 3x^2 - 5x - 1 = 0

myininaya (myininaya):

use the quadratic formula

OpenStudy (anonymous):

i know the formula i just confuse myself while applying it.

myininaya (myininaya):

a=,b=-5,c=-1 \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-(-5) \pm \sqrt{(-5)^2-4(3)(-1)}}{2(3)}\]

myininaya (myininaya):

\[x=\frac{5 \pm \sqrt{25+12}}{6}=\frac{5 \pm \sqrt{37}}{6}\]

OpenStudy (anonymous):

What do i do about the plus/minus symbol. idk how to use it

myininaya (myininaya):

you have solutions one with the + and one with the -

OpenStudy (anonymous):

thannk you:)

myininaya (myininaya):

\[\frac{5 \pm \sqrt{37}}{6}\] is the same as saying: \[\frac{5+\sqrt{37}}{6}. \frac{5-\sqrt{37}}{6}\]